# Exploring Parent’s Needs for Children-Centered AI to Support Preschoolers’ Storytelling and Reading Activities

### v1

@@ 4.4.4. Challenges in Supporting Situated and Adaptive Interaction
According to our participants, the practical storytelling was a dynamic and situated process. During the practical storytelling processes, parents frequently had a dynamic awareness of children’s real-time interests and attention levels, based on what to dynamically adjust the interaction contents. Furthermore, our participants expressed that even as parents, they may not have a comprehensive and correct understanding of their children’s characteristics and interests. Their understanding of their children and the interactive ways and contents they chose were gradually discovered and adjusted during the interactive process.
We also didn’t know what our child liked at the beginning. We can only adjust the interaction based on his feedback and state. For example, from his eyes and body language, we can learn that he doesn’t want to read anymore, he wants to go out to play, or he is not interested in this story. If we ask a question and he doesn’t react, we will change the question or the wording. Through this process, we gradually know what content and interactive methods he likes.” [P9]
Nonetheless, our participants stated that the majority of existing AI-based storytelling technologies were unable to dynamically perceive children’s real-time situations very well, which impacted the actual effectiveness of these tools directly. “Machines have a lot of knowledge and just want to impart it to children. They have no idea whether children want to listen or are interested in what they are saying.”, as P12 said.
@@ 4.4.5. Challenges in Supporting Children’s Active and Diversified Thinking
Throughout our study, our participants consistently emphasized that every child was unique. However, current technologies, they considered, fell short of adequately supporting the diverse development of children. For children-tailored conversational agents (e.g. C4 and C5 in Table 2), which were built on specialized children’s corpora and could ask children questions proactively, the interaction was primarily based on relatively fixed patterns and language models, which evaluated by our participants as “limiting children’s active and diversified development and thinking” [P11]. They thought children should engage actively in thinking and posing questions based on their curiosity, rather than relying on a machine to ask pre-determined questions and provide predefined answers. P4 explained this opinion through his using experience to C4 in Table 2:
Each story with it is followed by a universal truth, such as the stepmother is evil. But the child may be interested in stepmother’s beautiful dress, and its color. The process of proactive thinking and questioning is valuable. Even though many of the child’s questions may appear strange, silly, or stupid, they are actually their own questions and expressions of their interests.” [P4]
For general intelligent conversational agents based on massive corpora, such as C6 in Table 2, and even more intelligent generative AI technologies such as ChatGPT, many of our participants, such as P4 and P7, still expressed their rejection of using it to support children’s storytelling and interaction, with concerns of “standard answers” given by ChatGPT limiting the diverse development of children. They thought that, no matter how intelligent, AI technologies were fundamentally built upon extensive data-driven pattern recognition, i.e. mining rules based on massive data and providing ‘standardized’ and ‘correct’ answers, which were definitely not suitable for preschoolers, and even potentially mold them into fixed paradigms.
I don’t support using ChatGPT to tell stories or interact with children. ChatGPT can answer many questions, hundreds of thousands of questions, but these questions and their answers have already been summarized by predecessors. It standardizes these things and gives them to children, which actually limits children’s diversified thinking and development.
@@ 4.4.6. Challenges in Supporting Authenticity, Civilized, Safe and Ethical Interaction
Challenges in supporting authenticity, civilized, safe, and ethical interaction primarily pertained to general AI technologies such as C6 in Table 2, and in particular, the booming generative AI technologies due to their lack of transparency and insufficient controllability of LLM-based technologies (Wu et al., 2022). Our participants expressed that preschoolers did not have the ability to discern multiple information. Given the insufficient controllability, they were worried about what kind of conversations children would have with such a smart robot, and who could ensure the authenticity of the output content.
The value orientation of the content output by ChatGPT was also another concern expressed by our participants. P2 was a Christian and she explained this ”value orientation” to us: “In our family, we prioritize this aspect as we have strong beliefs. We focus on concepts that align with our faith, and scientific knowledge that contradicts our beliefs is not widely accepted. For example, we adhere to the creation theory rather than the evolution theory. I am unsure whether ChatGPT has the capability to accommodate or respect these specific preferences”[P2].
@@ 5. Discussion
This paper has illustrated how parents with preschoolers use, experience, and perceive the existing AI-based storytelling technologies and tools in practice. Our findings indicated that, while parents perceived some value when using these tools, they expressed that existing AI-based storytelling tools they had bought faced numerous practical challenges in supporting their expected storytelling. We categorize these challenges into interactive and algorithmic and elaborate on them in the Finding section. According to these findings, we now delve into parents’ expectations for ideal parent-child interactive storytelling and propose research implications for future AI-based storytelling research.
@@
5.1. Re-Conceptualizing the Interaction in AI-Mediated Storytelling Scenarios: Taking Parents as “Beneficiary”
@@
Most existing AI-based storytelling and related systems primarily focus on children-AI interaction. However, our research has revealed several challenges with this children-AI interaction paradigm. Specifically, preschoolers typically lack the suitability to use tools for storytelling independently and consistently, making solo storybook reading and telling with AI a challenging task. Meanwhile, the capabilities of existing AI tools are insufficient to meet parents’ expected storytelling. Furthermore, the process of storytelling for preschoolers is inherently a parent-children interactive experience, with one primary goal of enhancing parent-child relationships, as reported by our participants as well as existing literature (Wright, 1995; Chang and Huang, 2016; Peck, 1989). However, this aspect is not well-supported under existing Children-AI interaction paradigm.
@@
In CSCW, researchers are increasingly acknowledging the limitations of this Children-AI interaction paradigm (Zhang et al., 2022; Harbig et al., 2011), and also emphasizing the positive impact of parental involvement in children’s use of AI for storytelling (Cagiltay et al., 2022; Zhao and McEwen, 2021). Our study, echoing this literature, further emphasizes the importance of recognizing parents as direct beneficiaries of AI tools, and supporting an interactive paradigm among parents, children, and AI tools, rather than focusing solely on children-AI interaction. Based on our findings, We specifically propose the following design implications.
@@
Taking Parents as Beneficiaries: Support, not Substitute. We argue that the role of AI tools should be to assist parents in storytelling, rather than substituting their role entirely. According to our participants, despite the limitations of existing AI tools, they expressed expectations for AI tools to assist them by, for instance, alleviating some of the physical labor involved (e.g., helping them read stories), generating valuable and high-quality interactive content or questions for them when they lack expertise, providing inspiration for interactions with their children, and preparing appropriate resource for supporting their storytelling with children. These supporting functions could be considered by future AI tools.
@@
Taking Parents as Beneficiaries: Beyond a Storytelling Toolman. By taking parents as beneficiaries, we also suggest shifting parents’ role in AI-based storytelling from merely tool man or technology supporters, as most existing tools do, to active participants. Activities such as role-playing and interactive play can be incorporated. While role-playing is already utilized in some tools (Zhao and McEwen, 2021; Dietz et al., 2021), many of them still take parents as supporters in terms of helping with media content selection and text-to-speech usage, rather than fully participating as players. We recommend treating parents as true players to enhance their involvement. Ongoing research on interactive technologies (e.g., (Chen et al., 2022)) may offer potential support for better parent-child interactive storytelling experiences.
@@
Enabling Parents’ Personal Experience Through Co-Creating.. According to most of our participants, during their actual storytelling with preschoolers, their conversations often go beyond the storybooks, and integrate a lot of personal viewpoints and experiences, which has also been recognized in the literature as an essential and valuable way of storytelling (Vasalou et al., 2020; Zhang et al., 2022). We therefore propose AI-mediated storytelling could empower parents by involving them as co-storytellers, and providing them interfaces to seamlessly incorporate their personal experiences, childhood memories, or whatever they want into the narrative. Such integration could also effectively contribute to the realization of personalized story themes that align with parental objectives.
@@
In terms of interactive modality, our study has demonstrated that preschoolers often face certain cognitive difficulties in accepting the virtual interactive modality; Meanwhile, current AI tools often fail to effectively support preschoolers’ proactive and consistent participation. The main reason for these issues is that the current designs do not take into account the preschoolers’ cognitive and psychological features, resulting in a misalignment between child cognitive development and the virtual interactive way (Xu and Warschauer, 2020b).
According to these findings, we propose that future AI storytelling tools need to fit with preschoolers’ psychology and cognitive characteristics. Some classical theories of child cognitive development, such as Piaget’s theory of cognitive development (Piaget, 1976), could be used to guide the interactive design tailored for preschoolers. Within the framework of Piaget’s theory, each stage of development has distinct cognitive characteristics. Preschoolers are in the preoperational stage, where their cognitive processes remain anchored in concrete experiences, starting to be symbolic thinking. They can use imagination and symbols to express and solve problems, but their grasp of logical reasoning and abstract concepts is still limited(Piaget, 1976).
@@ 5.3. Design Considerations for Child-Appropriated AI Algorithms: Carefully Considering ”Appropriated”
In terms of core technologies and algorithms of existing AI-based storytelling tools, our participants reported significant challenges, primarily regarding the silly, unintelligent algorithms, and their failure to output adaptive and understandable content to children and answer their ”a hundred thousand whys”. From the technical perspective, a natural improving suggestion is to construct child-appropriated datasets and corpora that cater to preschoolers’ language and cognitive abilities, and then use them to enhance algorithm performance, making the output better suited to preschoolers’ needs in terms of language style, content, complexity, etc. Ongoing efforts have started to focus on these aspects, such as constructing question-answer storytelling datasets tailored to children’s needs (Xu et al., 2022; Chen et al., 2023), utilizing Knowledge Graph-based approaches for knowledge expansion (Chen et al., 2023), and improving AI to generate story book related, adaptive conversation (Yao et al., 2022; Kory and Breazeal, 2014). We believe that these research directions are crucial for the future AI-based storytelling technologies.
@@
Moving forward, we further suggest carefully considering what is ”appropriate” for preschoolers and, based on that, exploring directions for algorithmic improvement. When we refer to ”children-appropriated,” we are considering whether the interaction and generated content are suitable for children. This entails aligning with their cognitive level, personal abilities, interests, preferences, and other relevant factors (Moyer et al., 1987). However, from the algorithmic perspective, the ”appropriate” is often reflected in the technical performance, e.g. the accuracy of answers, which the majority of our participants believe may not be applicable for preschoolers, because there is no ”standard or correct answer” for each child. For instance, algorithms may focus on whether it can generate the correct answer “poisoned apple” for the question “what did Snow White eat and die from?”, while children may be more interested in questions like “What dress was the witch wearing?” We therefore suggest future algorithms to support more open and diverse interactions that encourage children’s creative and critical thinking, rather than solely focusing on providing the ”accurate” answer. The advancing large language models have the potential to support such conversational needs, and researchers have also started to examine its potential in supporting children’s creative expression and storytelling (Han and Cai, 2023a, b). We believe it is a valuable direction, but need to carefully consider content safety and ethical issues which are significant concerns raised by our participants.
@@
To conclude our discussion, we emphasize that not all challenges raised by parents can or should be resolved through AI algorithms, because, given AI’s understanding of ”appropriate”, it might be insufficient to answer the countless ”why” questions from preschoolers that have no correct answer, and support parents expected storytelling and personalized interactions. Further, AI solutions will inevitably involve the use of more perceptual technologies to gather additional data, posing challenges to the protection of children’s privacy. Our participation also expresses similar concerns in this regard. Echoing the discussion in section 5.1, we further emphasize it is crucial to carefully consider the role and functionality of AI tools in preschoolers’ storytelling - these tools should serve as aids for parents, enabling them to better support the storytelling process in a manner that aligns with their children’s unique needs.
@@ 5.4. Limitation and Future Work
Overall, our research presents an in-depth and empirical investigation of parents’ experiences and perceptions of AI-based storytelling. In particular, we elaborate on a range of challenges our participants reported in using existing AI-based tools to support their expected storytelling for preschoolers in practice. A major limitation of our study is the small sample size, which may result in possible sample selection bias (Heckman, 1979) and sample diversity (Wolda, 1981) issues. In particular, during the recruitment process, we did not consider the diverse characteristics of parents. However, during the research, we discovered that participants’ individual characteristics, such as their economic level, educational background, and professional expertise, significantly influenced their attitudes and perceptions toward using AI for storytelling. Unfortunately, given the sample size of our study, our current study did not analyze these perspectives very well. Furthermore, it is important to note that our study primarily took place in China. While this specific socio-technological context provided us with a unique opportunity to gain insights into AI-based storytelling practices, it also may limit the generalizability of our findings. In future research, we aim to expand our focus to a more diverse range of parental perspectives, painting a more complete picture of our research questions.
@@ 6. Conclusion
Using AI-based tools to support parent-children interactive storytelling has been a contemporary trend. In this work, we conducted an interview study to deeply examine how they use, experience, and perceive AI-based storytelling technologies in practical storytelling scenarios. Our findings suggest current AI-based storytelling technologies face numerous challenges in supporting storytelling for preschoolers as parents expect. We categorize them into interactive and algorithmic challenges, and elaborate on them in this paper. Drawing on these findings, we identified vital design implications to better facilitate the increasing demand for AI-based children-parent storytelling.
With the widespread applications of artificial intelligence in contemporary society, human-centered AI has become a focal point in CSCW, HCI and AI fields. However, how to effectively implement the human-centered vision and create practical technologies that truly put people first has consistently posed challenges. Our research focuses on the specific AI application scenario – AI-based storytelling for preschoolers, and delves into how the concept of human-centered AI can be realized in this scenario. While our research focuses on the specific scenario of storytelling, we believe that the design principles we discuss can be generalized to a broader range of technology designs aimed at facilitating children-AI interactions.
Acknowledgements.
To be added.

### v2

@@ 4.4.4. Challenges in Supporting Situated and Adaptive Interaction
According to our participants, the practical storytelling was a dynamic and situated process, in which parents had a dynamic awareness of children’s real-time interests and attention. Based on children’s feedback, parents dynamically adjusted the interaction content and ways. As P9 said,
We didn’t know what our child liked at the beginning. We adjust the interaction based on his feedback and state. For example, from his eyes and body language, we can learn that he doesn’t want to read anymore, he wants to go out to play, or he is not interested in this story. If we ask a question and he doesn’t react, we will change the question or the wording. Through this process, we gradually know what content and interactive methods he likes.” [P9]
Nonetheless, our participants stated that the majority of existing AI-based storytelling technologies were unable to perceive children’s real-time situations very well, which impacted the actual effectiveness of these tools directly. “Machines have a lot of knowledge and just want to impart it to children. They have no idea whether children want to learn or are interested in what they are saying.”, as P12 said.
@@ 4.4.5. Challenges in Supporting Children’s Active and Diversified Thinking
Throughout our study, our participants consistently emphasized that every child was unique. However, current technologies, they considered, fell short of adequately supporting the diverse development of children. For children-tailored conversational agents (e.g. C4 and C5 in Table 2), which were built on specialized children’s corpora and could ask children questions proactively, the interaction was primarily based on relatively fixed patterns and language models, which evaluated by our participants as “limiting children’s active thinking and diversified development” [P11]. They thought children should engage actively in thinking and posing questions based on their curiosity, rather than relying on a machine to ask pre-determined questions and provide predefined answers. P4 explained this opinion through his using experience to C4 in Table 2:
Each story with it is followed by a universal truth, such as the stepmother is evil. But the child may be interested in stepmother’s beautiful dress, and its color. The process of proactive thinking and questioning is valuable. Even though many of the child’s questions appear strange, silly, or stupid, they are actually their own questions and expressions of their interests.” [P4]
For general intelligent conversational agents based on massive corpora, such as C6 in Table 2, and even more intelligent generative AI technologies such as ChatGPT, many of our participants, such as P4 and P7, still expressed their rejection of using them, with concerns of “standard answers” limiting the diverse development of children. They thought that, no matter how intelligent, AI technologies were fundamentally built upon extensive data-driven pattern recognition, i.e. mining rules based on massive data and providing ‘standardized’ and ‘correct’ answers, which were definitely not suitable for preschoolers, and even potentially mold them into fixed paradigms.
I don’t support using ChatGPT to tell stories or interact with children. ChatGPT can answer many questions, hundreds of thousands of questions, but these questions and their answers have already been summarized by predecessors. It standardizes these things and gives them to children, which actually limits children’s diversified thinking and development.” [P4]
@@ 4.4.6. Challenges in Supporting Authenticity, Civilized, Safe and Ethical Interaction
Challenges in supporting authenticity, civilized, safe, and ethical interaction primarily pertained to general AI technologies such as C6 in Table 2, and the booming generative AI technologies due to their lack of transparency and insufficient controllability (Wu et al., 2022). Our participants expressed that preschoolers did not have the ability to discern multiple information. Given the insufficient controllability, they were worried about what kind of conversations children would have with such a smart robot, and who could ensure the authenticity of the output content.
The value orientation of the content output by ChatGPT was also another concern expressed by our participants. P2 was a Christian and explained this “value orientation” concern to us: “in our family, we prioritize this aspect as we have strong beliefs. We focus on concepts that align with our faith, and scientific knowledge that contradicts our beliefs is not widely accepted. For example, we adhere to the creation theory rather than the evolution theory. I am unsure whether ChatGPT has the capability to accommodate or respect these specific preferences”[P2].
@@ 5. Discussion
This paper has illustrated how parents with preschoolers in our study experience and perceive existing AI-based storytelling technologies and tools in practice. Our findings indicated that, despite some perceived value, existing AI-based storytelling tools our participants had bought faced numerous practical challenges in supporting their expected interactive storytelling. We categorize these challenges into interactive and algorithmic and elaborate on them. According to these findings, we now delve into parents’ expectations for ideal parent-child interactive storytelling and propose research implications for future AI-based storytelling research.
@@
5.1. Re-Conceptualizing Human-AI Interaction Paradigm in AI-based Storytelling Scenarios: Taking Parents as “Beneficiary”
@@
Most existing AI-based storytelling systems (UNICEF, 2021) primarily focus on child-AI interaction. However, our research has revealed several challenges with this child-AI interaction paradigm. Specifically, preschoolers typically lack the suitability to use tools for storytelling independently and consistently. Meanwhile, the capabilities of existing AI tools are insufficient to meet parents’ expected storytelling. Furthermore, the process of storytelling for preschoolers is inherently a parent-children interactive experience, with one primary goal of enhancing parent-child relationships, as reported by our participants as well as existing literature (Wright, 1995; Chang and Huang, 2016; Peck, 1989). However, this aspect is not well-supported under the existing child-AI interaction paradigm.
@@
In CSCW and related fields, researchers are increasingly acknowledging the limitations of this child-AI interaction paradigm (Zhang et al., 2022; Harbig et al., 2011), and also emphasizing the positive impact of parental involvement in children’s use of AI for storytelling (Cagiltay et al., 2022; Zhao and McEwen, 2021). Our study, echoing this literature, further emphasizes the importance of recognizing parents as direct beneficiaries of AI tools, and supporting an interactive paradigm among parents, children, and AI tools, rather than only focusing on child-AI interaction. Based on our findings, we specifically propose the following design implications.
@@
Taking Parents as Beneficiaries: Support, not Substitute. We argue that the role of AI tools should be assisting parents in storytelling, rather than substituting their role entirely. Despite the limitations of existing AI tools, our participants expressed expectations for AI tools to assist them by, for instance, alleviating some of the physical labor involved (e.g., helping them read stories), generating valuable and high-quality interactive content or questions for them when they lack expertise, providing inspiration for interactions with their children, and preparing appropriate resources for supporting their storytelling with children. These supporting functions could be considered by future AI-based interactive storytelling tools.
@@
Taking Parents as Beneficiaries: Beyond a Storytelling Toolman. With taking parents as beneficiaries, we also suggest shifting parents’ role in AI-based storytelling from merely tool man or technology supporters, as most existing tools do, to active participants. Activities such as role-playing and interactive play can be incorporated. While role-playing is already utilized in some tools (Zhao and McEwen, 2021; Dietz et al., 2021), many of them still take parents as supporters in terms of helping with media content selection and text-to-speech usage, rather than fully participating as players. We recommend treating parents as true players and enhancing their involvement. Ongoing research on interactive technologies (e.g., (Chen et al., 2022)) may offer potential support for better parent-child interactive storytelling experiences.
@@
Enabling Parents’ Personal Experience Through Co-Creating. According to most of our participants, during their actual storytelling with preschoolers, their conversations often go beyond the storybooks, and integrate a lot of personal viewpoints and experiences, which has also been recognized in the literature as an essential and valuable way of storytelling (Vasalou et al., 2020; Zhang et al., 2022). We therefore propose AI-based storytelling tools could empower parents by involving them as co-storytellers (i.e., human-AI collaboration (Wang et al., 2020)), and providing them interfaces to seamlessly incorporate their personal experiences, childhood memories, or whatever they want into the narrative. Such integration could also effectively contribute to the realization of personalized story themes that align with parental objectives.
@@
Our study has demonstrated that preschoolers often face certain cognitive difficulties in accepting the virtual interactive modality; Meanwhile, current AI tools often fail to effectively support preschoolers’ proactive and consistent participation. The main reason for these issues is that the current designs do not take into account the preschoolers’ cognitive and psychological features, resulting in a misalignment between child cognitive development and the virtual interactive way (Xu and Warschauer, 2020b).
According to these findings, we propose that future AI storytelling tools need to fit with preschoolers’ psychology and cognitive characteristics to make children be a more active role in storytelling processes. Some classical theories of child cognitive development, such as Piaget’s theory of cognitive development (Piaget, 1976), could be used to guide the interactive design tailored for preschoolers. Within the framework of Piaget’s theory, each stage of development has distinct cognitive characteristics. Preschoolers are in the preoperational stage, where their cognitive processes remain anchored in concrete experiences, starting to be symbolic thinking. They can use imagination and symbols to express and solve problems, but their grasp of logical reasoning and abstract concepts is still limited(Piaget, 1976).
@@ 5.3. Design Considerations for Child-Appropriated AI Algorithms: Carefully Considering ”Appropriated”
In terms of core technologies and algorithms of existing AI-based storytelling tools, our participants reported significant challenges, primarily regarding the silly, unintelligent algorithms, and their failure to output adaptive and understandable content to children and answer their ”a hundred thousand whys”. From the technical perspective, a natural improving suggestion is to construct child-appropriated datasets and corpora that cater to preschoolers’ language and cognitive abilities, and then use them to enhance algorithm performance, making the output better suited to preschoolers’ needs in terms of language style, content, complexity, etc. Ongoing efforts have started to focus on these aspects, such as constructing question-answer storytelling datasets tailored to children’s needs (Xu et al., 2022; Chen et al., 2023), utilizing Supervised Fine-Tuning (SFT (Xu et al., 2024)) or Knowledge Graph-Augmented Generation approaches (KG-RAG) for knowledge expansion (Chen et al., 2023), and improving AI to generate story book related, adaptive conversation (Yao et al., 2022; Kory and Breazeal, 2014). We believe that these research directions are crucial for the future AI-based storytelling technologies.
@@
Moving forward, we further suggest carefully considering what is ”appropriate” for preschoolers and, based on that, exploring directions for algorithmic improvement. When we refer to “children-appropriated”, we are considering whether the interaction and generated content are suitable for children. This entails aligning with their cognitive level, personal abilities, interests, preferences, and other relevant factors (Moyer et al., 1987). However, from the algorithmic perspective, the “appropriate” is often reflected in the technical performance, e.g. the accuracy of answers, which the majority of our participants believe may not be applicable for preschoolers, because there is no “standard or correct answer” for each child. For instance, algorithms may focus on whether it can generate the correct answer “poisoned apple” for the question “what did Snow White eat and die from?”, while children may be more interested in questions like “What dress was the witch wearing?” We therefore suggest future algorithms to support more open and diverse interactions that encourage children’s creative and critical thinking, rather than solely focusing on providing the ”accurate” answer. The advancing large language models have the potential to support such conversational needs, and researchers have also started to examine its potential in supporting children’s creative expression and storytelling (Han and Cai, 2023a, b). We believe it is a valuable direction, but need to carefully consider content safety and ethical issues which are significant concerns raised by our participants.
@@
To conclude our discussion, we emphasize that not all challenges raised by parents can or should be resolved through AI algorithms, because, given AI’s understanding of “appropriate”, it might be insufficient to answer the countless “why” questions from preschoolers that have no correct answer, and support parents expected storytelling and personalized interactions. Further, AI solutions will inevitably involve the use of more perceptual technologies to gather additional data, posing challenges to the protection of children’s privacy. Our participation also expresses similar concerns in this regard. Echoing the discussion in section 5.1, we further emphasize it is crucial to carefully consider the role and functionality of AI tools in preschoolers’ storytelling - these tools should serve as aids for parents, enabling them to better support the storytelling process in a manner that aligns with their children’s unique needs.
@@ 5.4. Limitation and Future Work
The major limitation of our study is the small sample size, which may result in possible sample selection bias (Heckman, 1979) and sample diversity (Wolda, 1981) issues. In particular, during the recruitment process, we did not consider the diverse characteristics of parents. However, during the research, we discovered that participants’ individual characteristics, such as their economic level, educational background, and professional expertise, significantly influenced their attitudes and perceptions toward using AI for storytelling. Yet, given the sample size of this study, we did not analyze these perspectives very well. Meanwhile, we recruited parents who had already inclined to use AI-based products. While such kind of participants could provide us with richer usage experiences with AI-based tools, the attitudes and perceptions of parents who have not used AI-based products might be ignored. In addition, all our participants were recruited from China, which might limit the generalizability of our findings.
In future research, we aim to expand our focus to a more diverse range of parental perspectives, painting a more complete picture of our research questions. Meanwhile, we will involve more stakeholders, including children and kindergarteners, to the study, and identify their practical needs and experiences, through which to shed light on future human-centered design in AI-based storytelling scenarios for preschoolers.
@@ 6. Conclusion
Using AI-based tools to support parent-child interactive storytelling has been a contemporary trend. In this work, we conducted an interview study to deeply examine how parents use, experience, and perceive AI-based storytelling technologies in practical storytelling scenarios. Our findings suggest current AI-based storytelling technologies face numerous challenges in supporting storytelling for preschoolers as parents expect. We categorize them into interactive and algorithmic challenges, and elaborate on them. Drawing on these findings, we identified vital design implications for AI technologies to better facilitate children-centered interactive storytelling.
With the widespread applications of AI in contemporary society, human-centered AI has become a focal point in CSCW, HCI and AI fields. However, how to effectively implement the human-centered vision and create practical technologies that truly put people first has consistently posed challenges. Our research focuses on the specific AI application scenario – AI-based interactive storytelling for preschoolers, and delves into how the concept of human-centered AI can be realized in this scenario. While our research focuses on the specific scenario of storytelling, we believe that the design principles we discuss can be generalized to a broader range of technology designs aiming at facilitating children-AI interaction.